# Deep Learning-based Intrusion Detection Systems: A Survey

### v1

@@ 1. Introduction
Generally, the intrusion detection process is automated through software or hardware systems, which are known as IDS (Rebecca and Mell, 2001; Liao et al., 2013; Nassar et al., 2019; Xue et al., 2021; Nasir et al., 2022; Ahmad et al., 2023; Awad et al., 2022; Goyal et al., 2023). Intrusion detection is the process of monitoring and analyzing the events occurring in a computer or a network for signs of intrusion. Intrusions occur when attackers access the systems via the Internet, as well as authorized users misuse their privileges or attempt to gain unauthorized privileges.
IDS play a critical role in protecting computer and network systems. The limitations of IDS may result in terrible damage to enterprises. One example is the recent Colonial Pipeline Ransomware Attack (Beerman et al., 2023). In April 2021, the hacking group DarkSide launched a ransomware attack on Colonial Pipeline, the biggest oil pipeline company in the United States, using an unused VPN account. Due to this attack, 5,500 miles of transportation pipelines were forced to shut down, affecting nearly 45% of the fuel supply on the Eastern Coast. The Colonial Pipeline paid \$4.4 million ransom money, in addition to the theft of over 100 GB of data. If the malware intrusion can be detected in time, the influence of this attack is allowed to be greatly mitigated or even eliminated.
@@ 1.1. Tough but Bright Intrusion Detection System
IDS have been increasingly challenged to effectively deal with intrusions for decades. It is noted in Figure 1(a) that the number of CVE [^1] Common Vulnerabilities and Exposures (CVE) is a security project for security information sharing and vulnerability management. CVE is a publicly accessible vulnerability recording database where each vulnerability has a common name and a unique identifier. records has presented an accelerating uptrend, especially in 2016, which suffered a sharp rise. After 2016, the number of CVE records stays growing at a high speed, reaching around 30,000 in 2024. Besides, according to the CNNVD [^2] Chinese National Vulnerability Database (CNNVD) is a Chinese national database that catalogs security vulnerabilities in software and hardware products. CNNVD also provides unique identifiers and descriptions similar to CVE. report shown in Figure 1(b), we can observe that almost all (i.e., 97.2%) vulnerabilities are medium risk or above, with high and critical risk accounting for 40% of them. The growing number of vulnerabilities and the large percentage of high-risk vulnerabilities both reveal the tough situation faced by IDS.
Nevertheless, an interesting observation from Figure 1(a) is that, against the number of CVE records, DL-IDS papers also started to emerge in 2016 and their amount grew year by year subsequently. We can notably find that the growth trend of DL-IDS papers is nearly the same as that of CVE records. The potential reason can be speculated as DL is an effective way for IDS to cope with their tough situation. Borrowing the strong generalizability from DL techniques, DL-IDS detection can be extended to zero-day intrusions that are almost impossible to detect with the traditional DL-IDS. Some studies (Yang et al., 2023b; Wang et al., 2022a; Zengy et al., 2022) demonstrate this speculation. In their experiments, DL-IDS are all reported with an achievement of over 90% detection accuracy while the traditional DL-IDS sometimes only have around 50% detection accuracy.
@@ 1.2. Related Surveys and Our Scope
Unfortunately, none of the related surveys in the last decade have systematically investigated DL-IDS. On one hand, some related surveys may only focus on a few parts of DL-IDS, such as log parsers (S and Nasreen, 2021; Lupton et al., 2021; Zhang et al., 2023), datasets (Stojanović et al., 2020) and attack modeling (Stojanović et al., 2020; Alshamrani et al., 2019). On the other hand, while several surveys (He et al., 2020; Inam et al., 2023; Lv et al., 2022; Li et al., 2021; Zipperle et al., 2022; Nassar et al., 2019; Martins et al., 2022; Bridges et al., 2019; Nasir et al., 2022; Khraisat et al., 2019; Liao et al., 2013) involve some DL-based approaches, they did not review DL-IDS from the perspective of DL particularly.
@@
The surveys (Zhang et al., 2023; S and Nasreen, 2021; Lupton et al., 2021; Stojanović et al., 2020; Alshamrani et al., 2019) are the typical example papers describing only a few parts of DL-IDS. Among them, Adel (Alshamrani et al., 2019) mainly studied various techniques and solutions that were tailored to APT attacks, as well as discussed where to make the APT detection framework smart. Scott et al. (Lupton et al., 2021) and Tejaswini et al. (S and Nasreen, 2021) detailedly discussed online log parsers and their applications for anomaly detection. Branka et al. (Stojanović et al., 2020) review APT datasets and their creation, along with feature engineering in attack modeling. Zhang et al. (Zhang et al., 2023) created an exhaustive taxonomy of system log parsers and empirically analyzed the critical performance and operational features of 17 open-source log parsers. For DL-IDS, all the above surveys are obviously insufficient to advance research understanding and provide theoretical suggestions.
@@
Another type of existing surveys engaged DL-IDS but studied them from the other perspectives (He et al., 2020; Inam et al., 2023; Lv et al., 2022; Li et al., 2021; Zipperle et al., 2022; Nassar et al., 2019; Martins et al., 2022; Bridges et al., 2019; Nasir et al., 2022; Khraisat et al., 2019; Liao et al., 2013; Ahmad et al., 2021). Specifically, the surveys (Khraisat et al., 2019; Liao et al., 2013) aim to give an elaborate image of IDS and comprehensively explain methods from signature checking to anomaly detection algorithms. Originating from log data, the survey (He et al., 2020) presented a detailed overview of automated log analysis for reliability engineering and introduced three tasks including anomaly detection, failure prediction, and failure diagnosis. In survey (Nasir et al., 2022), Nasir et al. explored the efficacy of swarm intelligence on IDS and highlighted the corresponding challenges in multi-objective IDS problems.
@@
Our work distinguishes from the related surveys by providing a comprehensive literature review of DL-IDS. From the perspective of DL, our survey elaborates a common workflow of DL-IDS and introduces the corresponding taxonomies of all modules within this workflow. Moreover, our survey discusses the possible challenges and research visions for DL-IDS, which include many DL-related issues that have not yet been studied by the existing surveys.
@@ 1.3. Contributions and Organization
  • Realizing that IDS has made significant progress with the aid of DL, we present a thorough survey for DL-IDS, formalizing its definition and clarifying its type location.
  • We outline the common workflow for DL-IDS, consisting of the data management stage and intrusion detection stage. Further, we systematically illustrate the research advances in all modules of this workflow and innovatively taxonomize the papers based on DL techniques.
@@
Section 2 introduces the survey methodology of this work. Section 3 describes the background knowledge about DL-IDS. Section 4 and Section 5 elaborates the recent research trends on data management and intrusion detection, respectively. Section 6 illustrates the benchmark datasets and their feature dimensions. Section 7 discusses the visions and challenges for future research. Lastly, the conclusion is presented in Section 8.
@@ 2. Survey Methodology
To start our literature review, we selected several popular literature databases, including Web of Science (Analytics, 1997), IEEE Xplore (IEEE, 2000), and Scopus (Elsevier, 2021), as the search engine. For search keywords, we determined from generalized terms associated with DL-IDS, such as intrusion detection system, attack investigation, anomaly detection, threat detection, Advanced Persistent Threats, data provenance analysis, forensic analysis, causality analysis, log collection, log compression, log parsing, log storage, and log summarization. Then, we employed Connected Papers (Papers, 2020), a visual tool that assists researchers in finding relevant academic papers, to ensure that we did not overlook the typical related literature. Since the found literature is numerous and rather generalized for the DL-IDS scope, we carefully checked their topics and prioritized only academic papers that are significantly related to DL. Finally, all these papers were filtered based on the impact factors of their published journals or academic conferences, leaving us a total of 42 papers.
We identified a few venues that have published many significant papers in the field of DL-IDS. They can be broadly divided into three categories: network security, software and system, and database. We report the distribution of these papers with their published years in Figure 2.
@@ 3.1.2. Types of IDS
Figure 3. Types of IDS.
Generally, IDS can be further categorized into various types based on their data source. Well-known types include NIDS (Meng et al., 2020; Min et al., 2021; Trivedi et al., 2023; Qazi et al., 2023), HIDS (Xie et al., 2018; Wang et al., 2018; Han et al., 2020; Wang et al., 2020; Shin and Kim, 2020; Wei et al., 2021a; Fält et al., 2021; Anjum et al., 2022; Wu et al., 2022; Zengy et al., 2022; Wang et al., 2022a; Yang et al., 2023b; Jia et al., 2024; Rehman et al., 2024; Kapoor et al., 2021; Liao et al., 2023), and Provenance-based IDS (PIDS) (Xie et al., 2018; Han et al., 2020; Wang et al., 2020; Wei et al., 2021a; Anjum et al., 2022; Wu et al., 2022; Zengy et al., 2022; Wang et al., 2022a; Yang et al., 2023b; Jia et al., 2024; Rehman et al., 2024; Kapoor et al., 2021; Liao et al., 2023). In Figure 3, we present the taxonomy of these IDS types and their corresponding data sources, along with the location of DL-IDS within the taxonomy.
@@
NIDS takes network traffic between hosts as its input. It is usually deployed at the edge or key node of the network, allowing it to secure the whole computer system with limited data. Benefiting from the global perception of the whole computer system, NIDS does well in large-scale multi-host intrusions such as Distributed Denial-of-Service (DDoS) attacks. However, NIDS performs poorly in intra-host intrusions and is unable to analyze intrusions in the form of encrypted network traffic.
@@
PIDS is a subtype of HIDS, particularly referring to HIDS that utilizes data provenance as its data source. Due to analysis in the intact trail of events, PIDS is proven to be effective in coping with advanced attacks (Zipperle et al., 2022). By performing causality analysis on data provenance, PIDS can significantly reduce false alarms compared with traditional IDS. However, data provenance can be very expensive to obtain, requiring complicated technical tools for monitoring operating systems, network protocols, and applications.
@@
Unlike the other types of IDS such as NIDS and HIDS are categorized by their data sources, DL-IDS is defined by the techniques used in intrusion detection. As shown in Figure 3, the data source of DL-IDS can be network traffic, system events, or both. Taking advantage of the generalizability of DL techniques, DL-IDS is allowed to handle zero-day attacks precisely and thus become extremely interested in the cybersecurity community recently. In this work, we comprehensively review DL-IDS, detailing its entire workflow and describing its visions and challenges.
@@ 3.2. Common Workflow
Figure 4 illustrates the common workflow of DL-IDS. It usually consists of 7 steps: raw data, collection, storage, parsing, summarization, detection, and investigation. The brief introductions of these steps are as follows:
@@
  • Storage involves managing large amounts of collected log data, with storage engines and blockchain methods offering solutions for reliable and efficient performance.
@@
Note that DL-IDS can also be performed in other step orders by skipping some of the steps. For example, log data can be first parsed before storage (Liu et al., 2019). Attack investigation can be directly conducted without detection of intrusions (Alsaheel et al., 2021). Basically, the above steps of DL-IDS can be divided into two stages: data management stage and intrusion detection stage.
In the data preprocessing stage, security analysts collect crucial data to describe the states of computer systems via various data collection tools. This data is then efficiently and reliably stored and managed with the aid of ELK and blockchain methods. Lastly, log parsers are applied to organize the data into structured ones in which only key information are reserved.
In the intrusion detection stage, data is utilized at the semantic level. Based on its semantics, redundant and missing values of data are merged or removed to significantly reduce the computing costs. Subsequently, DL-based detection methods are developed to detect anomalies within data to identify the presence of intrusions. In the end, by conducting a causality analysis of the anomalies, intrusion scenarios can be recovered as detection reports that are provided to security analysts.
Our survey will be structured according to these two stages, illustrating the data management stage in Section 4 and the intrusion forensics stage in Section 5.
@@ 4.1. Data Collection
The first step of DL-IDS is to collect useful data from raw data. Raw data indicates records that document events, activities, and operations that occur within a system, application, or network (a.k.a., logs), represented by network traffic or audit logs. Generally, network traffic refers to data transmitted over a network, governed by standardized network protocols. And audit logs mainly refer to chronological record data generated by systems or applications. By collecting useful logs, DL-IDS is allowed to monitor the health condition and operational status of information systems (Zhang et al., 2023). Common attributes of logs include timestamp, event type, subject, object, description, etc.
However, on different platforms, logs possess different formats and organizational structures (Zhang et al., 2023; Zipperle et al., 2022; Bridges et al., 2019; Li et al., 2021). To counter this, researchers have created various log collection tools specialized for various systems. For example, in Windows systems, Event Viewer is employed to manage system logs. Yet in Linux systems, log files are usually saved in the /var/log/ directory. The classification of data collection tools is shown in Table 1, including Windows, Linux, Cloud, and Cross platforms.
@@ 4.1.1. Windows Platform Tools
Event Tracing for Windows (ETW) (Microsoft, [n. d.]) is a powerful event tracing mechanism provided by Microsoft. It consists of three components: providers, controllers, and consumers. ETW instrument applications to provide kernel event logging and allows developers to start and stop event tracing sessions momentarily.
Panorama (Yin et al., 2007) exploits hardware-level and OS-aware dynamic taint tracking to collect logs. Moreover, it develops a series of automated tests to detect malware based on several kinds of anomalous behaviors.
@@ 4.1.2. Linux Platform Tools
auditd (Grubb, 2008) is a native collection tools supported by the Linux kernel, which is responsible for writing audit logs to disk and monitoring a variety of auditable events such as system calls, file accesses, and modifications.
sysdig (Kili, [n. d.]) relies on the kernel module to achieve monitoring and data collection of the system. sysdig focuses on system runtime monitoring and fault troubleshooting, which is also widely used in containers and cloud-native environments.
CamFlow (Pasquier et al., 2017) designs a self-contained, easily maintainable implementation of whole-system provenance based on Linux Security Module, NetFilter, and other kernel facilities. Furthermore, it provides a mechanism to adapt the captured data provenance to applications and can be integrated across distributed systems.
Tracee (Tracee, 2022) takes advantage of the extended Berkeley Packet Filter (eBPF) framework to observe systems efficiently. It uses eBPF to tap into systems and expose that information as events.
DataTracker (Stamatogiannakis et al., 2015) is an open-source data provenance collection tool using dynamic instrumentation. It is able to identify data provenance relations of unmodified binaries without access to or knowledge of the source codes.
Inspector (Thalheim et al., 2016) is a Portable Operating System Interface (POSIX)-compliant data provenance library for shared-memory multi-threaded applications. It is implemented as a parallel provenance algorithm on a concurrent provenance graph.
AutoLog (Huo et al., 2023) generates runtime log sequences by analyzing source codes and does not need to execute any programs. It can efficiently produce log datasets (e.g., over 10,000 messages/min on Java projects) and has the flexibility to adapt to several scenarios.
eAudit (Sekar et al., 2024) is a scalable and easily deployable data collection tools. eAudit relies on the eBPF framework built into recent Linux versions, making it work out of the box on most of the Linux distributions.

### v2

@@ 1. Introduction
The automated system for detecting intrusions is known as IDS. The limitations of IDS may result in terrible damage to enterprises. One example is the recent Colonial Pipeline Ransomware Attack (Beerman et al., 2023). In April 2021, the hacking group DarkSide launched a ransomware attack on Colonial Pipeline, the biggest oil pipeline company in the United States, using an unused VPN account. Due to this attack, 5,500 miles of transportation pipelines were forced to shut down, affecting nearly 45% of the fuel supply on the Eastern Coast. The Colonial Pipeline paid \$4.4 million ransom money, in addition to the theft of over 100 GB of data. If the malware intrusion can be detected in time, the influence of this attack can be greatly mitigated or even eliminated.
@@ 1.1. Tough but Bright Intrusion Detection System
IDS have been increasingly challenged to effectively deal with intrusions for decades. It is noted in Figure 1(a) that the number of CVE [^1] Common Vulnerabilities and Exposures (CVE) is a security project for security information sharing and vulnerability management. CVE is a publicly accessible database where each vulnerability has a common name and a unique identifier. records has presented an accelerating uptrend, especially in 2016, which suffered a sharp rise. After 2016, the number of CVE records stays growing at a high speed, reaching around 30,000 in 2024. Besides, according to the CNNVD [^2] Chinese National Vulnerability Database (CNNVD) is a Chinese national database that catalogs security vulnerabilities in software and hardware products. CNNVD also provides unique identifiers and descriptions similar to CVE. report shown in Figure 1(b), we can observe that almost all (i.e., 97.2%) vulnerabilities are medium risk or above, with high and critical risk accounting for 40% of them. The growing number of vulnerabilities and the large percentage of high-risk vulnerabilities both reveal the tough situation faced by IDS.
Nevertheless, an interesting observation from Figure 1(a) is that, against the number of CVE records, DL-IDS papers also started to emerge in 2016 and their amount grew year by year subsequently. We can notably find that the growth trend of DL-IDS papers is nearly the same as that of CVE records. The potential reason can be speculated as DL is an effective way for IDS to cope with their tough situation. Borrowing the strong generalizability from DL techniques, DL-IDS detection can be extended to zero-day intrusions that are almost impossible to detect with the traditional DL-IDS. Some studies (Yang et al., 2023; Wang et al., 2022; Zengy et al., 2022) demonstrate this speculation. In their experiments, DL-IDS are all reported with an achievement of over 90% detection accuracy while the traditional DL-IDS sometimes only have around 50% detection accuracy.
@@ 1.2. Related Surveys and Our Scope
Unfortunately, none of the related surveys in the last decade have systematically investigated DL-IDS. On one hand, some related surveys may only focus on a few parts of DL-IDS, such as log parsers (S and Nasreen, 2021; Lupton et al., 2021; Zhang et al., 2023b), datasets (Stojanović et al., 2020), attack modeling (Stojanović et al., 2020; Alshamrani et al., 2019), and specific DL technique type (Bilot et al., 2023). On the other hand, while several surveys (He et al., 2020; Inam et al., 2023; Lv et al., 2022; Li et al., 2021; Zipperle et al., 2022; Nassar et al., 2019; Martins et al., 2022; Bridges et al., 2019; Nasir et al., 2022; Khraisat et al., 2019; Liao et al., 2013) involve some DL-based approaches, they did not review DL-IDS from the perspective of DL particularly.
@@
The surveys (Zhang et al., 2023b; S and Nasreen, 2021; Lupton et al., 2021; Stojanović et al., 2020; Alshamrani et al., 2019) are the typical example papers describing only a few parts of DL-IDS. Among them, Adel et al. (Alshamrani et al., 2019) mainly studied various techniques and solutions that were tailored to APT attacks, as well as discussed where to make the APT detection framework smart. Scott et al. (Lupton et al., 2021) and Tejaswini et al. (S and Nasreen, 2021) detailedly discussed online log parsers and their applications for anomaly detection. Branka et al. (Stojanović et al., 2020) review APT datasets and their creation, along with feature engineering in attack modeling. Zhang et al. (Zhang et al., 2023b) created an exhaustive taxonomy of system log parsers and empirically analyzed the critical performance and operational features of 17 open-source log parsers. Tristan et al. (Bilot et al., 2023) focused on the applications of graph neural networks (GNNs) to IDS. For DL-IDS, all the above surveys are obviously insufficient to advance research understanding and provide theoretical suggestions.
@@
Another type of existing surveys involved DL-IDS but studied them from the other perspectives (He et al., 2020; Inam et al., 2023; Lv et al., 2022; Li et al., 2021; Zipperle et al., 2022; Nassar et al., 2019; Martins et al., 2022; Bridges et al., 2019; Nasir et al., 2022; Khraisat et al., 2019; Liao et al., 2013; Ahmad et al., 2021). Specifically, the surveys (Khraisat et al., 2019; Liao et al., 2013) aim to give an elaborate image of IDS and comprehensively explain methods from signature checking to anomaly detection algorithms. Originating from log data, the survey (He et al., 2020) presented a detailed overview of automated log analysis for reliability engineering and introduced three tasks including anomaly detection, failure prediction, and failure diagnosis. In survey (Nasir et al., 2022), Nasir et al. explored the efficacy of swarm intelligence on IDS and highlighted the corresponding challenges in multi-objective IDS problems.
@@
Our work distinguishes the related surveys by providing a comprehensive literature review of DL-IDS. From the perspective of DL, our survey elaborates on a common workflow of DL-IDS and introduces the corresponding taxonomies of all modules within this workflow. Moreover, our survey discusses the possible challenges and research visions for DL-IDS, which include many DL-related issues that have not yet been studied by the existing surveys.
@@ 1.3. Contributions and Organization
  • Realizing that IDS has made significant progress with the aid of DL over the last decade, we present a thorough survey for DL-IDS, formalizing its definition and clarifying its location among other types of IDS.
  • We outline the common workflow for DL-IDS, consisting of the data management stage and intrusion detection stage. We further systematically illustrate the research advances in all modules of this workflow and innovatively taxonomize the papers based on DL techniques.
@@
Section 2 introduces the survey methodology of this work. Section 3 describes the background knowledge about DL-IDS. Section 4 and Section 5 elaborates the recent research trends on data management stage and intrusion detection stage, respectively. Section 6 illustrates the benchmark datasets and their feature dimensions. Section 7 discusses the visions and challenges for future research. Lastly, the conclusion is presented in Section 8.
@@ 2. Survey Methodology
Figure 3. Types of IDS.
To start our literature review, we selected several popular literature databases, including Web of Science (Analytics, 1997), IEEE Xplore (IEEE, 2000), and Scopus (Elsevier, 2021), as the search engine. For search keywords, we determined from generalized terms associated with DL-IDS, such as intrusion detection system, attack investigation, anomaly detection, threat detection, Advanced Persistent Threats, data provenance analysis, forensic analysis, causality analysis, log collection, log compression, log parsing, log storage, and log summarization. Then, we employed Connected Papers (Papers, 2020), a visual tool that assists researchers in finding relevant academic papers, to ensure that we did not overlook the typical related literature. Since the found literature is numerous and rather generalized for the DL-IDS scope, we carefully checked their topics and prioritized only academic papers that are related to DL. Finally, all these papers were filtered based on the impact factors of their published journals or academic conferences, leaving us a total of 113 papers.
We identified a few venues that have published many significant papers in the field of DL-IDS, such as Security, S&P, CCS, NDSS, TIFS, TDSC, ICSE, ASE, ESEC/FSE, TSE, EuroSys, NSDI, OSDI, SOSP, ATC, ICDE, TKDE, SIGMOD, KDD, WWW, and IJCAI. We broadly divide them into four categories: security, software, system, and data. The distribution of these papers with their published years is reported in Figure 3.
@@ 3.1.2. Types of IDS
Generally, IDS can be further categorized into various types based on their data sources (Zipperle et al., 2022). Well-known types include NIDS, HIDS, and Provenance-based IDS (PIDS). Figure 3 depicts IDS types, their data sources, and the location of DL-IDS within those IDS types.
@@
NIDS takes network traffic between hosts as its input. It is usually deployed at the edge or key node of the network, allowing it to secure the whole computer system with limited data. Benefiting from the global perception of the whole computer system, NIDS does well in large-scale multi-host intrusions such as Distributed Denial-of-Service (DDoS) attacks. However, NIDS performs poorly in intra-host intrusions and is difficult to analyze intrusions in the form of encrypted network traffic.
@@
PIDS is a subtype of HIDS, particularly referring to HIDS that utilizes data provenance as its data source. Due to analysis in the intact trail of events, PIDS is proven effective in coping with advanced attacks (Zipperle et al., 2022). By performing causality analysis on data provenance, PIDS can significantly reduce false alarms. Yet, data provenance is very expensive to obtain, requiring complicated technical tools for monitoring operating systems, network protocols, and applications.
@@
Unlike the other types of IDS such as NIDS and HIDS are categorized by their data sources, DL-IDS is defined by the techniques used in intrusion detection. As shown in Figure 3, the data source of DL-IDS can be network traffic, system events, or both. Taking advantage of the generalizability of DL techniques, DL-IDS is allowed to handle zero-day attacks precisely and thus become extremely interested in the cybersecurity community recently.
@@ 3.2. Common Workflow
Figure 4 depicts the common workflow of DL-IDS. It usually consists of 7 steps: raw data, collection, storage, parsing, summarization, detection, and investigation, which are explained as follows:
@@
  • Storage involves storage and search engines to manage large amounts of collected log data. Log data is labeled with indexes for efficient retrieval.
@@
Note that DL-IDS can also be performed in other step orders by skipping some of the steps. For example, log data can be first parsed before storage (Liu et al., 2019). Attack investigation can be directly conducted without detection of intrusions (Alsaheel et al., 2021). This survey is organized by the common workflow.
@@ 4.1. Data Collection
The first step of DL-IDS is to collect useful data from raw data. Raw data indicates records that document events, activities, and operations that occur within a system, application, or network (a.k.a., logs), represented by audit logs or application logs within hosts, or network traffic between hosts. By collecting useful logs, DL-IDS is allowed to monitor the health condition and operational status of information systems (Zhang et al., 2023b). Common attributes of logs include timestamp, event type, subject, object, description, etc.
On different platforms, logs possess different formats and organizational structures (Zhang et al., 2023b; Zipperle et al., 2022; Bridges et al., 2019; Li et al., 2021). To counter this, researchers have created various log collection tools specialized for various systems. For example, in Windows systems, Event Viewer is employed to manage system logs. Yet in Linux systems, log files are usually saved in the /var/log/ directory. The classification of data collection tools is shown in Table 1, including Windows, Linux, Cloud, and Cross platforms.
@@
ISDC (Mirnajafizadeh et al., 2024) Eliminating overheads on in-network resources
@@ 4.1.1. Windows Platform Tools
Event Tracing for Windows (ETW) (Microsoft, [n. d.]) is a powerful event tracing mechanism provided by Microsoft. It consists of three components: providers, controllers, and consumers. ETW instruments applications to provide kernel event logging and allows developers to start and stop event tracing sessions momentarily. Panorama (Yin et al., 2007) exploits hardware-level and OS-aware dynamic taint tracking to collect logs. Moreover, it develops a series of automated tests to detect malware based on several kinds of anomalous behaviors.
@@ 4.1.2. Linux Platform Tools
auditd (Grubb, 2008) is a native collection tool supported by the Linux kernel, which is responsible for writing audit logs to disk and monitoring a variety of auditable events such as system calls, file accesses, and modifications. sysdig (Kili, [n. d.]) relies on the kernel module to achieve monitoring and data collection of the system. sysdig focuses on system runtime monitoring and fault troubleshooting, which is also widely used in containers and cloud-native environments. CamFlow (Pasquier et al., 2017) designs a self-contained, easily maintainable implementation of whole-system provenance based on Linux Security Module, NetFilter, and other kernel facilities. Furthermore, it provides a mechanism to adapt the captured data provenance to applications and can be integrated across distributed systems. Tracee (Tracee, 2022) takes advantage of the extended Berkeley Packet Filter (eBPF) framework to observe systems efficiently. It uses eBPF to tap into systems and expose that information as events. DataTracker (Stamatogiannakis et al., 2015) is an open-source data provenance collection tool using dynamic instrumentation. It is able to identify data provenance relations of unmodified binaries without access to or knowledge of the source codes. Inspector (Thalheim et al., 2016) is a Portable Operating System Interface (POSIX)-compliant data provenance library for shared-memory multi-threaded applications. It is implemented as a parallel provenance algorithm on a concurrent provenance graph. AutoLog (Huo et al., 2023) generates runtime log sequences by analyzing source codes and does not need to execute any programs. It can efficiently produce log datasets (e.g., over 10,000 messages/min on Java projects) and has the flexibility to adapt to several scenarios. eAudit (Sekar et al., 2024) is a scalable and easily deployable data collection tools. eAudit relies on the eBPF framework built into recent Linux versions, making it work out of the box on most of the Linux distributions.